OpenStudy (sushi121212):

Write a new equation: f(x)=-x+5; horizontal translation 2 units left

satellite73 (satellite73):

find \[f(x+2\) to move it two units left

OpenStudy (mathmale):

Begin with the given equation: f(x)=-x+5 Replace "x" in this equation with "( )" (on both sides) write "x+2" inside each of the 2 sets of parentheses Simplify. Done.

OpenStudy (sushi121212):

f(x+2)=(x+2)+5 f(x+2)=x+7 Did I do that right?

OpenStudy (danjs):

Almost, you forgot the negative on the x. f(x+2)= -(x + 2) + 5 = -x + 3

OpenStudy (sushi121212):

Oh, okay, thank you!!